OpenStudy (anonymous):

Wendy’s pay is $423.00. 19% of that is subtracted for taxes. How much is left after paying taxes?

OpenStudy (texaschic101):

423 - 423(.19) = 423 - 80.37 = $ 342.63

OpenStudy (mathmale):

texaschic101 has done a nice job here. First, she wrote out the "gross pay," $423. Then, she calculated the 19% tax, $80.37. Lastly, she subtracted the tax from the gross pay to obtain the net pay (paycheck amount).